# Quality Diversity through Human Feedback

### v2

@@ 5 Related Work
This work expands upon recent developments in methodologies for aligning models with human objectives, especially reinforcement learning from human feedback (RLHF) (Christiano et al., 2017; Ibarz et al., 2018). RLHF was initially proposed for training RL agents in simulated environments such as Atari games. It has been later applied to fine-tune or perform one-shot learning on language models for tasks such as text summarization (Ziegler et al., 2019; Stiennon et al., 2020; Wu et al., 2021), dialogue (Jaques et al., 2019), and question-answering (Nakano et al., 2021; Bai et al., 2022; Ouyang et al., 2022), as well as vision tasks such as perceptual similarity measure (Fu et al., 2023). While past work focuses on learning a reward or preference model from human intentions, we propose to learn diversity metrics through human feedback and use them to drive the optimization process in QD algorithms.
@@
Instead of optimizing for one optimal solution, diversity-driven optimization methods such as Novelty Search (Lehman & Stanley, 2011a) and Quality Diversity (Lehman & Stanley, 2011b; Cully et al., 2015; Mouret & Clune, 2015; Pugh et al., 2016) aim to identify a variety of (top-performing) solutions that exhibit novelty or diversity. Prior work expands on QD by enhancing diversity maintenance (Fontaine et al., 2019; Smith et al., 2016; Vassiliades et al., 2017), search process (Fontaine et al., 2020; Vassiliades & Mouret, 2018; Nordmoen et al., 2018; Sfikas et al., 2021), and optimization mechanism (Kent & Branke, 2020; Conti et al., 2018; Fontaine & Nikolaidis, 2021). This work focuses on situations where, unlike the general assumption in these methods, there is no predefined diversity metrics. Recent work (Grillotti & Cully, 2022; Cully, 2019) explores unsupervised methods for diversity discovery, but the resulting diversity often fail to support optimization. Our work differs by leveraging human feedback to derive diversity that are aligned with human interest, and thus often more beneficial for optimization.
@@ 6 Conclusion
This paper introduced Quality Diversity through Human Feedback (QDHF), which expands the reach of QD algorithms through leveraging human feedback to effectively derive diversity metrics. Empirical results show that QDHF outperforms current unsupervised diversity discovery methods for QD, and compares well to standard QD that utilizes manually-crafted diversity metrics. In particular, applying QDHF in a latent space illumination task shows that QDHF substantially enhances the diversity of images in the text-to-image generation task. We provide analysis of QDHF’s scalability and the quality of diversity metrics it learns. Future work will focus on applying QDHF in more challenging robotics, RL, and generative tasks in more complex, open-ended environments.
@@ Acknowledgments
Li Ding and Lee Spector were supported by the National Science Foundation under Grant No. 2117377. Jeff Clune and Jenny Zhang were supported by the Vector Institute, a grant from Schmidt Futures, an NSERC Discovery Grant, and a generous donation from Rafael Cosman. The authors would like to thank members of the PUSH lab at Amherst College, University of Massachusetts Amherst, and Hampshire College for helpful discussions, and anonymous reviewers for their thoughtful feedback. This work was performed in part using the high performance computing equipment from Collaborative R&D Fund managed by Massachusetts Technology Collaborative.
@@ Appendix A Preliminaries on Quality Diversity
A.1 Quality Diversity
The QD process aims to find the best representative samples, not only seeking the absolute best but also ensuring that the selections are varied and uniquely excellent in their own ways. Intuitively, imagine assembling a soccer team with QD: it meticulously recruits top-tier players across various positions to build a well-rounded team, rather than simply gathering the most renowned players regardless of their specialized roles. Key aspects of Quality Diversity include:
  • Exploration and Exploitation: QD balances exploration (searching for new, diverse solutions) and exploitation (refining known good solutions). This balance helps in thoroughly understanding the solution space and uncovering unique solutions that might be overlooked by conventional methods.
A.2 MAP-Elites
@@ Appendix B Additional Implementation Details
For the maze navigation task, the solution is the network parameters of the default MLP policy network with a hidden-layer size of 8. The episode length of the environment is 250. We evaluate the policy and obtain the state descriptors. The objective is the accumulated reward at each state. For diversity measures, the ground truth diversity is the end-position of the agent, i.e., the position at the last state. For QDHF and AURORA, we extract features from the state descriptor as the x and y positions of the agent at each state. The latent projection in QDHF transforms the feature into a 2-dim embedding. For AURORA, the auto-encoder has the same architecture of 64-32-2-32-64 nodes in each layer, where the mid-layer is the embedding used for QD. For QDHF, we use 200 judgments of simulated human feedback. For all experiments, we run Map-Elites for 1000 iterations, and for each iteration, we generate a batch of 200 solutions with Gaussian mutation (adding Gaussian noises sampled from $\mathcal{N}(0,0.2^{2})$), and evaluate them. The archive has a shape of $(50,50)$.
@@
Latent space illumination.
In the LSI task, we run QDHF for 200 iterations with a batch size of 5 solutions per iteration. The solutions are generated with Gaussian mutation (adding Gaussian noises sampled from $\mathcal{N}(0,0.1^{2})$). The archive has a shape of $(20,20)$. The solution is the latent vector used as the input to Stable Diffusion, which has a shape of $(4,64,64)$. We use Stable Diffusion v2.1-base, which generates images at a resolution of 512x512. The feature extractor is a CLIP model with ViT-B/16 backbone, which returns a 512-dim feature vector. QDHF learns a latent projection from 512-d to 2-d. To gather online human feedback, we use DreamSim with the DINO-ViT-B/16 backbone. The DreamSim model is trained on the NIGHTS dataset, which consists of 20k synthetic image triplets annotated with human judgments as labels. For QDHF, we use 10000 judgments of predicted human feedback. The objective is the CLIP score of the image and the text prompt. The text prompt is also input to the Stable Diffusion model to condition the generation towards more relevant content.
Appendix C Additional Experimental Results
C.1 Robotic Arm
Alignment between learned and ground truth diversity metrics.
Table 4: List of objective prompts for the LSI task.
prompt #1 a photo of an astronaut riding a horse on mars
prompt #2 an image of a bear in a national park
prompt #3 an image of a cat on the sofa
prompt #4 an image of a person playing guitar
prompt #5 an image of a dog in the park
prompt #6 an image of urban downtown
Table 5: Detailed quantitative results for the LSI task.
Best-of-N Best-of-N+ QDHF
# CLIP MeanPD Std.PD CLIP MeanPD Std.PD CLIP MeanPD Std.PD
1 $72.15$ $0.422$ $0.083$ $\mathbf{72.19}$ $0.421$ $0.082$ $71.85$ $\mathbf{0.565}$ $\mathbf{0.152}$
2 $69.79$ $0.293$ $0.109$ $69.78$ $0.293$ $0.107$ $\mathbf{69.94}$ $\mathbf{0.434}$ $\mathbf{0.144}$
3 $69.19$ $0.418$ $0.094$ $69.16$ $0.423$ $0.093$ $\mathbf{69.47}$ $\mathbf{0.541}$ $\mathbf{0.155}$
4 $\mathbf{68.35}$ $0.437$ $0.120$ $68.33$ $0.431$ $0.119$ $67.95$ $\mathbf{0.571}$ $\mathbf{0.147}$
5 $68.90$ $0.564$ $0.117$ $68.97$ $0.569$ $0.119$ $\mathbf{69.43}$ $\mathbf{0.603}$ $\mathbf{0.167}$
6 $64.74$ $0.388$ $0.113$ $65.01$ $0.382$ $0.113$ $\mathbf{65.83}$ $\mathbf{0.448}$ $\mathbf{0.140}$
Figure 5: Summary of user’s levels of experience with text-to-image geneartion tools.
Table 6: Detailed user experience study results for the LSI task. The survey data is collected on 46 participants through an online form.
User Preference User Perceived Diversity
Best-of-N QDHF Hard to Decide Best-of-N QDHF Hard to Decide
1 $\mathbf{22}$ $19$ $5$ $2$ $\mathbf{43}$ $1$
2 $12$ $\mathbf{24}$ $10$ $2$ $\mathbf{36}$ $8$
3 $8$ $\mathbf{30}$ $8$ $3$ $\mathbf{38}$ $5$
4 $14$ $\mathbf{29}$ $3$ $8$ $\mathbf{36}$ $2$
5 $7$ $\mathbf{30}$ $9$ $5$ $\mathbf{25}$ $16$
6 $10$ $\mathbf{20}$ $16$ $4$ $\mathbf{31}$ $11$
@@ C.2 Latent Space Illumination
We include more detailed results on the LSI task with different prompts. Table 4 lists the prompts that we used as objectives for the LSI task.
Quantitative evaluation.
Table 5 includes detailed quantitative results for each prompt. We can see that QDHF consistently demonstrates better diversity than the baselines measured by pairwise distance using DreamSim, and shows comparable generation quality as measured by CLIP score, which is the objective for optimization.
User experience study.
Table 6 includes detailed user experience study results for each prompt, which summarizes data from a survey of 46 participants with varying experience with text-to-image software. We ask the user’s experience with text-to-image generation tools in the beginning of the survey. The summary is visualized in Fig. 5. We conduct a “blind” study, where participants are unaware of the origin of each output when making the decision. A sample question used in the user feedback survey is shown in Fig. 6.
@@
We also show examples of generated images in Fig. 7 to Fig 12. For all the figures, the top 4x4 grid displays images with the highest CLIP scores from randomly generated images. The bottom grid displays a uniformly-sampled subset of QDHF solutions. Qualitatively, images generated by QDHF have more variations, and show visible trends of diversity.
@@
Notably, while QDHF significantly outperforms the baseline on most cases, we find that prompt #5 and #6 are two sub-performing cases for QDHF. For both prompts, the diversity of QDHF results is not apparent. The most likely reason is that the preference model (DreamSim) does not generalize well to cases such as different breeds of dogs, scenes of a park, and appearances of cityscapes. We aim to solve the above issues in future work where human feedback needs to be collected in a more diverse and strategic way to facilitate better generalization of the preference model, and thus improves the performance of QDHF.
@@
Another interesting finding is that for prompt #1, while most users find QDHF results are more diverse, more than half of the users actually prefers the less diverse baseline results. According to the feedback from users, people may prefer less diverse but more content-focused results in some specific cases. The relationship between diversity and user preference under different use cases in generative AI application remains an open question, and we look forward to exploring this topic in future work.
Figure 6: An example of questions in the user feedback survey.
@@
Figure 7: Qualitative results for prompt #1: “a photo of an astronaut riding a horse on mars”. The top 4x4 grid shows the baseline results and the bottom grid shows QDHF results.
Figure 8: Qualitative results for prompt #2: “an image of a bear in a national park”. The top 4x4 grid shows the baseline results and the bottom grid shows QDHF results.
Figure 9: Qualitative results for prompt #3: “an image of a cat on the sofa”. The top 4x4 grid shows the baseline results and the bottom grid shows QDHF results.
@@
Figure 10: Qualitative results for prompt #4: “an image of a person playing guitar”. The top 4x4 grid shows the baseline results and the bottom grid shows QDHF results.
Figure 11: Qualitative results for prompt #5: “an image of a dog in the park”. The top 4x4 grid shows the baseline results and the bottom grid shows QDHF results.
@@
Figure 12: Qualitative results for prompt #6: “an image of urban downtown”. The top 4x4 grid shows the baseline results and the bottom grid shows QDHF results.
kjXdWy329+dTgeSJD3ieZ7RNO0VAXAPwDEAO5VKndi2fWrb9jWl9Esul6PZbDY9Go1OZ7PZ9z/lyuD3OozU2wAAAABJRU5ErkJggg==" alt="[LOGO]">

### v3

@@ 6 Related Work
This work expands upon recent developments in methodologies for aligning models to human objectives, especially reinforcement learning from human feedback (RLHF) (Christiano et al., 2017; Ibarz et al., 2018). RLHF was initially proposed to train RL agents in simulated environments such as Atari games. It has since been applied to fine-tune or perform one-shot learning on language models for tasks including text summarization (Ziegler et al., 2019; Stiennon et al., 2020; Wu et al., 2021), dialogue (Jaques et al., 2019), and question-answering (Nakano et al., 2021; Bai et al., 2022; Ouyang et al., 2022), as well as vision tasks such as measuring perceptual similarity (Fu et al., 2023). While past efforts have focused on learning reward or preference models from human intentions, we propose to learn diversity metrics through human feedback, which then drives the optimization process in QD algorithms.
Figure 5: Visualization of maze navigation solutions in different diversity spaces. “GT” stands for ground truth. Each point on the heatmap is a solution with its objective value visualized in color. QDHF fills up the archives with more solutions than AURORA. While both methods learned a rotated version of the maze as the diversity metrics (first column), QDHF more accurately captures the scale of the maze, especially in under-explored areas.
@@
Instead of optimizing for a single optimal solution, diversity-driven optimization methods such as QD (Lehman & Stanley, 2011b; Cully et al., 2015; Mouret & Clune, 2015; Pugh et al., 2016) and Novelty Search (Lehman & Stanley, 2011a) aim to identify a variety of (top-performing) solutions that exhibit novelty or diversity. Prior work expands on QD by enhancing diversity maintenance (Fontaine et al., 2019; Smith et al., 2016; Vassiliades et al., 2017), the search process (Fontaine et al., 2020; Vassiliades & Mouret, 2018; Nordmoen et al., 2018; Sfikas et al., 2021), optimization mechanism (Kent & Branke, 2020; Conti et al., 2018; Fontaine & Nikolaidis, 2021), and exploring implicit methods for diversity-driven optimization (Ding & Spector, 2022; Boldi et al., 2023; Ding et al., 2023; Ni et al., 2024; Spector et al., 2024). Recent work (Grillotti & Cully, 2022a, b; Cully, 2019) also explores unsupervised methods for diversity discovery. Our work differs by leveraging human feedback to derive diversity metrics that are aligned with human interest, and thus often more beneficial for optimization.
@@ 7 Conclusion
This paper introduces Quality Diversity through Human Feedback (QDHF), which expands the reach of QD by leveraging human feedback to effectively derive diversity metrics. Empirical results show that QDHF outperforms current unsupervised diversity discovery methods, and compares well to standard QD that uses manually crafted diversity metrics. In particular, applying QDHF in an open-ended generative task substantially enhances the diversity of text-to-image generations. We provide an analysis of QDHF’s scalability, robustness, and the quality of diversity metrics learned. Future work will focus on applying QDHF to more challenging tasks in complex and open-ended environments.
@@ Acknowledgments
Li Ding and Lee Spector were supported by the National Science Foundation under Grant No. 2117377. Jeff Clune and Jenny Zhang were supported by the Vector Institute, a grant from Schmidt Futures, an NSERC Discovery Grant, and a generous donation from Rafael Cosman. Any opinions, findings, and conclusions or recommendations expressed in this publication are those of the authors and do not necessarily reflect the views of the funding agencies. The authors would like to thank Andrew Dai and Herbie Bradley for insightful suggestions, Bryon Tjanaka for help preparing the tutorial, members of the PUSH lab at Amherst College, University of Massachusetts Amherst, and Hampshire College for helpful discussions, and anonymous reviewers for their thoughtful feedback. This work utilized resources from Unity, a collaborative, multi-institutional high-performance computing cluster at the Massachusetts Green High Performance Computing Center (MGHPCC).
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ Appendix A Preliminaries on Quality Diversity
A.1 Quality Diversity
The QD process aims to find the best representative samples, not only seeking the absolute best but also ensuring that the selections are varied and uniquely excellent in their own ways. Intuitively, imagine assembling a soccer team with QD: it meticulously recruits top-tier players across various positions to build a well-rounded team, rather than simply gathering the most renowned players regardless of their specialized roles. Key aspects of QD include:
  • Exploration and Exploitation: QD balances exploration (searching for new, diverse solutions) and exploitation (refining known good solutions). This balance helps navigate the solution space and uncover unique solutions that may be overlooked by conventional methods.
A.2 MAP-Elites
@@ Appendix B Additional Implementation Details
For all tasks.
The frequency of diversity representation fine-tuning in QDHF decreases exponentially over time as the learned metrics become more robust. In our experiments, the latent projection is updated 4 times at iteration $1$, $10\%\cdot n$, $25\%\cdot n$, and $50\%\cdot n$ for a total of $n$ iterations. To fairly compare with QDHF-base, each update consumes 1/4 of the total budget of human feedback.
@@
For the maze navigation task, the robot features two contact sensors, and each of these sensors yields a value of 1 upon contact with a wall and -1 in the absence of any contact. The robot operates within a bounded square maze comprised of immovable, straight walls, which prevents the robot from moving upon collision. The episode length of the environment is 250. The solution is the network parameters of the default MLP policy network with a hidden-layer size of 8. We evaluate the policy and obtain the state descriptors. The objective is the accumulated reward at each state. For diversity measures, the ground truth diversity is the end-position of the agent, i.e., the position at the last state. For QDHF and AURORA, we extract features from the state descriptor as the x and y positions of the agent at each state. The latent projection in QDHF transforms the feature into a 2-dim embedding. For AURORA, the auto-encoder has the same architecture of 64-32-2-32-64 nodes in each layer, where the mid-layer is the embedding used for QD. For QDHF, we use 200 judgments of simulated human feedback. For all experiments, we run MAP-Elites for 1000 iterations, and for each iteration, we generate a batch of 200 solutions with Gaussian mutation (adding Gaussian noises sampled from $\mathcal{N}(0,0.2^{2})$), and evaluate them. The archive has a shape of $(50,50)$.
Latent space illumination.
In the LSI task, we use human judgment data from the NIGHTS dataset, and the DreamSim model to estimate image similarity, both from Fu et al. (2023). For comparisons, we implement a best-of-n approach on Stable Diffusion as the baseline, which generates a total of 2,000 images (the same number generated by QDHF) with latents randomly sampled from uniform $\mathcal{U}(0,1)$, and select a solution set of 400 images (the same number of solutions to be produced by QDHF) with the top CLIP scores. We also propose a stronger baseline featuring a heuristic (Best-of-n+), which works as follows: Starting from the second image to be generated, we sample 100 latents randomly and choose the one with a maximal l2 distance to the previously sampled latent. This approach increases the diversity of the latents and could potentially also increase the diversity of responses.
We run QDHF for 200 iterations with a batch size of 5 solutions per iteration. The solutions are generated with Gaussian mutation (adding Gaussian noises sampled from $\mathcal{N}(0,0.1^{2})$). The archive has a shape of $(20,20)$. The solution is the latent vector used as the input to Stable Diffusion, which has a shape of $(4,64,64)$. We use Stable Diffusion v2.1-base, which generates images at a resolution of 512x512. The feature extractor is a CLIP model with ViT-B/16 backbone, which returns a 512-dim feature vector. QDHF learns a latent projection from 512-d to 2-d. To gather online human feedback, we use DreamSim with the DINO-ViT-B/16 backbone. The DreamSim model is trained on the NIGHTS dataset, which consists of 20k synthetic image triplets annotated with human judgments as labels. For QDHF, we use 10000 judgments of predicted human feedback. The objective is the CLIP score of the image and the text prompt. The text prompt is also input to the Stable Diffusion model to condition the generation towards more relevant content.
Appendix C Additional Results
In this section, we include more experimental results to for further analysis of the proposed method.
C.1 Robotic Arm
Alignment between learned and ground truth diversity.
@@ C.1 Robotic Arm
Figure 7: Example of the human evaluation interface on Amazon Mechanical Turk.
Table 5: Results for robotic arm with noisy labels. We report the QD score (normalized to a scale of 0-100) and coverage for “All Solutions” (solutions found throughout training) and “Archive Solutions” (solutions in the final archive). We additionally report the performance of QDHF at different noise levels, where 5% Noise means 5% of the preference labels being randomly sampled are flipped. QDHF’s performance is not significantly affected at the low noise level (5%), and still outperforms QDHF-Base (with perfect data) and AURORA at the highest noise level of 20%.
All Solutions Archive Solutions
Methods QD Score Coverage QD Score Coverage
AURORA-Pre (AE) $38.5\pm 12.7$ $56.2\pm 6.8$ $14.3\pm 6.3$ $20.3\pm 6.4$
AURORA-Inc (AE) $53.0\pm 9.4$ $63.2\pm 6.2$ $17.6\pm 4.1$ $18.9\pm 4.1$
AURORA-Pre (PCA) $38.4\pm 13.6$ $54.1\pm 9.0$ $14.2\pm 4.4$ $19.3\pm 5.5$
AURORA-Inc (PCA) $45.9\pm 6.4$ $59.0\pm 3.7$ $18.3\pm 3.4$ $19.0\pm 3.5$
QDHF-Base $54.5\pm 4.3$ $62.7\pm 2.7$ $31.8\pm 4.5$ $34.1\pm 4.2$
QDHF (5% Noise) $72.4\pm 0.8$ $76.9\pm 1.0$ $55.6\pm 0.7$ $58.7\pm 0.5$
QDHF (10% Noise) $68.3\pm 4.7$ $73.8\pm 4.2$ $44.5\pm 11.5$ $47.2\pm 12.0$
QDHF (20% Noise) $62.6\pm 5.1$ $70.6\pm 2.4$ $35.4\pm 10.1$ $37.9\pm 9.9$
QD-GT $74.8\pm 0.2$ $79.5\pm 0.3$ $74.8\pm 0.2$ $79.5\pm 0.3$
Noisy feedback.
To investigate the impact of noise in human feedback data on the performance of QDHF, we conducted an ablation study on the robotic arm task. In this study, we manually injected noise into the feedback data, which was simulated from a ground truth diversity metric. We selected three noise levels—5%, 10%, and 20%—and at each level, a corresponding percentage of the feedback data, randomly sampled from the entire dataset, was flipped. This simulates potential bias or errors by human annotators during the annotation process.
The results, as presented in Table 5, indicate that QDHF’s performance is not significantly affected at the lower noise level of 5%. Although QDHF’s performance declines at higher noise levels, it still outperforms the baseline method (with perfect data) and all variants of AURORA, even at the highest noise level of 20%. These findings validate QDHF’s robustness against potential noise in the feedback data from humans.
@@ C.2 Latent Space Illumination
We include more detailed experimental setups and results for the LSI task.
Quantitative results on compositional prompts.
We include the quantitative evaluations of results on compositional prompts in Table 6, which are summarized over 100 text prompts. QDHF has a similar CLIP score to SD (Best-of-n), but much higher mean and standard deviation of pairwise distance, which indicates that QDHF is able to generate more diverse solutions while maintaining the high-quality. We can see that even though the CLIP scores is high, the issue of attribute leakage or interchanged attributes still exists in the results. However, QDHF is able to mitigate this issue by exploring more diverse responses, which results in a higher chance of approaching the correct response. For a more accurate evaluation of correctness, we rely on human evaluation as shown in Table 4.
Table 6: Quantitative Results for LSI (compositional prompts). We report the CLIP score and DreamSim pairwise distance (mean and std.) as quantitative metrics. QDHF demonstrates competitive quality to Stable Diffusion (Best-of-n) measured by CLIP score, and significantly better diversity measured by pariwise distance.
Method CLIP Score Mean PD Std. PD
SD (Best-of-n) $71.58$ $0.410$ $0.088$
QDHF $\mathbf{71.77}$ $\mathbf{0.543}$ $\mathbf{0.128}$
User study details.
The LSI experiments are conducted with singular prompts and compositional prompts. We use human evaluation to assess the results through a “blind” user study, where the users are unaware of the origin of each output when making the decision. For singular prompts, we use 100 prompts and each prompt produces results of two 4x4 grids of images for SD (Best-of-N) and QDHF, respectively. We evaluate the results by asking the users to answer two questions: (1) which set of generated images is more preferred, and (2) which set of generated images is more diverse. The study is conducted on Amazon Mechanical Turk, where each evaluation task is created as a Human Intelligence Task (HIT). For each prompt, we seek 50 evaluations from different users, and a user is qualified if they have a HIT Approval Rate over $95\%$. We follow the Amazon MTurk guidelines to fairly compensate the annotators for \$0.01 per evaluation per text prompt. An example of the MTurk annotation interface is shown in Fig. 7.
For compositional prompts, we follow a similar human evaluation procedure, and the question is: which set of generated images has the most accurate response. We also conduct experiments on 100 text prompts, sampled from the MCC-250 benchmark (Bellagente et al., 2023), with 50 evaluations per prompt. The comparisons are done on 3x3 grids of images since the evaluation target is on correctness rather than diversity.
More qualitative results on singular prompts.
@@
We show more qualitative results of generated images with singular prompts in Fig. 8 to Fig 11. For all the figures, the left 4x4 grid displays images with the highest CLIP scores from randomly generated images, and the right grid displays a uniformly sampled subset of QDHF solutions. Qualitatively, images generated by QDHF have more variations and show visible trends of diversity.
@@
Notably, while QDHF significantly outperforms the baseline on most cases, we find that Fig 11 is a sub-performing case for QDHF. Although QDHF is able to generate both day and night scenes, the diversity between the scenes is not apparent. The most likely reason is that the preference model (DreamSim) does not generalize well to cases such as different appearances of cityscapes. We aim to solve the above issues in future work where human feedback needs to be collected in a more diverse and strategic way to facilitate better generalization of the preference model and thus improve the performance of QDHF.
@@
Another interesting finding is that for prompt Fig 8, while most users find QDHF results are more diverse, more than half of the users actually prefer the less diverse baseline results. According to the feedback from users, people may prefer less diverse but more content-focused results in some specific cases. The relationship between diversity and user preference under different use cases in generative AI applications remains an open question, and we look forward to exploring this topic in future work.
@@
Figure 8: Qualitative results for singular prompt: “a photo of an astronaut riding a horse on mars”.
Figure 9: Qualitative results for singular prompt: “an image of a cat on the sofa”.
@@
Figure 10: Qualitative results for singular prompt: “an image of a person playing guitar”.
@@
Figure 11: Qualitative results for singular prompt: “an image of urban downtown”.
More qualitative results on compositional prompts.
We also include more qualitative results of generated images with compositional prompts in Fig. 12 and Fig 13. For all the figures, the left 3x3 grid displays images with the highest CLIP scores from randomly generated images, and the right grid displays a uniformly sampled subset of QDHF solutions. Qualitatively, images generated by QDHF have more variations and show a higher frequency of correctness.
Figure 12: Qualitative results for compositional prompt: “a red boat and a blue car”.
Figure 13: Qualitative results for compositional prompt: “a blue cake and a gold clock”.